OpenStudy (anonymous):

integrate (e^tanx)/(cos^2x)dx from 0 to pie/4

OpenStudy (anonymous):

hi jackie! :)

OpenStudy (anonymous):

heya

OpenStudy (anonymous):

\[\large \int\limits\limits \frac {e^{tanx}}{\cos^2x} dx\]

OpenStudy (anonymous):

is it that?

OpenStudy (anonymous):

oh.. wait...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

from 0 to pi/4

OpenStudy (anonymous):

\[\huge \int\limits\limits\limits \frac {e^{tanx}}{\cos^2x} dx\] is it that?

OpenStudy (anonymous):

u-sub... u=tanx

OpenStudy (anonymous):

ook

OpenStudy (anonymous):

want a hint to make it easy? write it as \[\int_0^{\frac{\pi}{4}}e^{\tan(x)}sec^2(x)dx\] then a u-sub suggests itself right away

OpenStudy (anonymous):

ok... there u go...

OpenStudy (anonymous):

yeah i guess so. but when it comes to taking finals i probably wont remember to do it that way.

OpenStudy (anonymous):

when handling integrals involving trig functions, consider the identities you can use to simplify things... if possible...:)

OpenStudy (anonymous):

ok? got an answer?

OpenStudy (anonymous):

no, confused.